OpenStudy (anonymous):

Find the equation of a line w/a slope of 1/2 passing through (-4, 2). Do I use y=mx+b?

OpenStudy (anonymous):

slope =1/2 point=(-4,2)

OpenStudy (anonymous):

yess u use y?mx+b just put x and y on equation...

OpenStudy (anonymous):

use point slope

OpenStudy (anonymous):

(y-y1)=m(x-x1)

OpenStudy (anonymous):

and then find b... 2 = 1/2 * -4 + b

OpenStudy (anonymous):

point slope is easier. Then you don't have to solve for b and re plug it in, you can solve it in one step.

OpenStudy (anonymous):

ahh malevolence is right ^^